### v1

# An Investigation of Incorporating Mamba for Speech Enhancement

Abstract
This work aims to study a scalable state-space model (SSM), Mamba, for the speech enhancement (SE) task. We exploit a Mamba-based regression model to characterize speech signals and build an SE system upon Mamba, termed SEMamba. We explore the properties of Mamba by integrating it as the core model in both basic and advanced SE systems, along with utilizing signal-level distances as well as metric-oriented loss functions. SEMamba demonstrates promising results and attains a PESQ score of 3.55 on the VoiceBank-DEMAND dataset. When combined with the perceptual contrast stretching technique, the proposed SEMamba yields a new state-of-the-art PESQ score of 3.69.
Index Terms—  Mamba, speech enhancement, state-space machine, consistency loss, time sequence modeling
Fig. 1: Architecture of our basic Mamba-based Speech Enhancement (SE) model, SEMamba-basic.
1 Introduction
A speech enhancement (SE) process involves retrieving clean speech components from a distorted signal to generate an improved signal with enhanced acoustic properties  [1, 2]. In various speech applications, such as assistive hearing technologies  [3, 4], speaker recognition  [5, 6], and automatic speech recognition  [7, 8], SE serves as a critical front-end processor. Recently, SE has been framed as a regression task utilizing neural-network-based mapping functions to convert noisy speech signals into cleaner outputs. More specifically, deep neural network models are employed to implement the mapping function for enhancing noisy speech. Several deep architectures, including deep denoising autoencoders [9], deep fully connected neural networks [10, 11], convolutional neural networks [12], long short-term memory  [13], convolutional recurrent neural networks [14, 15], and more recently, Transformer and its variant Conformer, have been used to form the mapping function, resulting in notably promising SE performance  [16, 17, 18, 19].
In addition to exploring advanced model architectures, several studies aimed to develop effective objective functions with the goal of boosting the SE capability. Typically, a signal-level distance measure usually serves as the fundamental objective function for training, e.g., L1/L2 norm  [20], SI-SDR  [21], or multiple-resolution loss  [22]. Objective functions specifically crafted and optimized for the downstream SE tasks have also been proposed. For instance, objective scores, such as perceptual evaluation of speech quality (PESQ)  [23], or short-time objective intelligibility (STOI)  [24], are learned using neural network models in  [25, 26] and then employed to optimize the SE model. Similarly, in  [27], subjective assessment results are learned through a model and utilized for training the SE model. Furthermore, a few studies have shown improved SE results by preprocessing the target speech prior to model training. For example, perceptual contrast stretching (PCS)  [28] enhances the contrast of target features based on their perceptual importance, effectively boosting SE performance without a notable increase in computational cost during runtime.
Recently, a highly promising neural architecture, referred to as Mamba  [29], has emerged, harnessing the state-space model with a selection mechanism. Mamba exhibits comparable or superior performance to state-of-the-art (SOTA) Transformer-based models across diverse tasks, due to its intrinsic capability in modeling extremely long-range dependencies. Furthermore, Mamba stands out for its efficient use of computational resources.
This research presents an in-depth comparison of Transformer based and Mamba-based SE models and introduces a new Mamba-based SE system, named SEMamba. First, we evaluate SEMamba using a basic SE system, termed SEMamba-basic, and compare its performance against a Transformer-based counterpart. Subsequently, we implement SEMamba with an advanced architecture, termed SEMamba-advanced. Experimental results indicate that SEMamba-advanced attains PESQ scores of 3.55 without PCS and 3.69 with PCS on the VoiceBank-DEMAND dataset [30].
2 Mamba: Linear-Time Sequence Modeling with Selective State Spaces
The structured state space model (SSM) [31] has demonstrated the ability to manage long-dependent sequences with low computation and memory requirements. It can serve as a replacement for either a CNN for efficient parallel training or an RNN for rapid autoregressive generation  [32]. More recently, Mamba  [29] has introduced significant advancements in discrete data modeling of SSM by introducing two key enhancements. First, Mamba incorporates an input-dependent selection mechanism, enabling efficient information filtering from inputs by parameterizing the SSM parameters based on the input data. Second, Mamba introduces a hardware-aware algorithm that scales linearly with input sequence length, facilitating faster computation of the model recurrently with a scan. Moreover, the Mamba architecture, which integrates SSM blocks with linear layers, is notably simpler and has demonstrated SOTA performance across various long-sequence patterns, including language and genomics, highlighting significant computational efficiency during both training and inference phases.
Structured SSMs, as described in the Mamba method, operate by mapping an input x to an output y through a higher dimensional latent state h, as follows:
$h_{n}$ $=\bar{\textbf{{A}}}h_{n-1}+\bar{\textbf{{B}}}x_{n},$ (1)
$y_{n}$ $=\textbf{{C}}h_{n}$ (2)
where $\bar{\textbf{{A}}}$ and $\bar{\textbf{{B}}}$ represent discretized state matrices. The discretization process converts continuous parameters $(\bf{\Delta},\textbf{{A}},\textbf{{B}})$ into discrete counterparts $(\bar{\textbf{{A}}},\bar{\textbf{{B}}})$, allowing the model to process discrete-time audio signals  [29].
The Mamba architecture integrates components from the H3 architecture  [33] and a gated multilayer perceptron block into a stacked structure, thereby expanding the model dimension by a factor and focusing most parameters in linear projections. This design features fewer parameters for the inner SSM and adopts the SiLU/Swish activation function, along with standard normalization and residual connections. For further details, the interested reader is referred to  [29].
3 Mamba in Speech Enhancement
We implement two types of the SEMamba solution. In the first solution, we integrate Mamba with a basic SE model architecture, as shown in Fig. 1. The second solution follows the advanced MP-SENet architecture but using the Mamba block in place of attention-based methods, as shown in Fig. 2. More details are given in the following sections.
3.1 SEMamba-basic
We first explore the construction of SEMamba using a basic causal model architecture. In this causal framework, the output at time $t$, $y_{t}$, is influenced only by inputs $x_{n}$ with $n\in\{0,\dots,t\}$, which implies a causal SE system. As depicted in Fig. 1, the input noisy waveform is first transformed into its spectral representation using the Short-Time Fourier Transform (STFT). Next, the magnitude (Mag) component of the STFT is compressed using the $log1p$ function $(log1p(z)=log(1+z))$  [34] to regulate the dynamic range of the magnitude. The compressed spectral magnitude is then fed into the SE module, which comprises (i) a convolutional encoder having 4 convolutional layers, (ii) a uni-directional Mamba block, and (ii) a fully connected decoder layer. The enhanced spectral magnitude features are decompressed by an inverse operation of $log1p$ before being combined with the noisy phase in the iSTFT block to reconstruct the enhanced waveform.
For comparative analysis, we also implement another basic SE solution using the Transformer block. Specifically, the Mamba SE block in Fig. 1 is replaced with a Transformer block featuring eight self-attention heads. Comparing those two architectures allows us to evaluate the difference between the attention and the selective state mechanisms.
Fig. 2: Architecture of the proposed SEMamba-advanced with Time-Frequency (TF) and Selective-SSM mechanism.
3.2 SEMamba-advanced
In the advanced solution, the causality constraint is removed; moreover, a more advanced SE structure is used. Specifically, the SE block integrates components from MP-SENet  [18], which concurrently denoises magnitude and phase spectra, serving as the backbone of our architecture. In detail, a spectral representation is obtained from the input noisy waveform using the STFT. The magnitude component is first compressed and then stacked together with the phase components. These components are processed by a feature encoder, featuring a dilated DenseNet core flanked by two convolutional layers. The output of the feature encoder undergoes multiple transformations via the Time-Frequency Mamba block, which is executed $N$ times ($N$=4 in this study) to enhance its spectral properties. The output of the Time-Frequency Mamba block is then routed to two separate decoders for magnitude and phase processing, respectively. Each decoder comprises a dilated DenseNet followed by a deconvolutional block, and it ends with an output layer that employs a 2D-convolution. The output layers reconstruct the magnitude mask and the real and imaginary parts of the waveform, respectively.
The loss function used to train the neural paramters follows that proposed in MP-SENet. Therefore, it is a linear combination of losses, including PESQ-based GAN discriminator, time, magnitude, complex, and phase losses.
3.3 SEMamba-advanced & additional designs
3.3.1 From uni- to bi-directional Mamba
While the original causal Mamba module has shown effectiveness, we further explore its performance by transitioning from a uni-directional to a bi-directional Mamba. This modification involves processing the input sequences both in their original and flipped forms. Specifically, the inputs are processed in parallel through the Mamba, after which the outputs are concatenated. The combined output then undergoes a $Conv1D$ layer, namely:
$\textbf{y}=Conv1D(M_{uni}(x)\oplus M_{uni}(flip(\textbf{x}))),$
where x, y, $M_{uni}()$, $flip()$, $Conv1D()$, and $\oplus$, respectively, denote input, output, uni-directional Mamba, flipping operation, 1-D convolution, and concatenation.
3.3.2 Consistency loss (CL)
Stability during the training process can be improved by leveraging the consistency loss proposed in [35]. The consistency loss aims to minimize the gap between the complex spectrum obtained directly from the model output (comprising amplitude and phase) and the complex spectrum derived after applying inverse Short-Time Fourier Transform (iSTFT) and then re-applying STFT to the resulting waveform. This gap arises because the complex spectrum predicted directly by the model may fall outside the STFT domain. Once processed through iSTFT and then subjected to STFT, the spectrum re-enters the STFT domain, thereby closing this gap. This mechanism ensures that our enhancements in the complex domain translate effectively in the time-frequency domain.
3.3.3 Perceptual contrast stretching (PCS)
PCS is a spectral processing technique that aims to improve the perceptual quality of speech signals. It leverages empirical observations of varying sensitivity levels in the human auditory system. PCS exploits this phenomenon by stretching the magnitude spectrum of the signal based on each frequency band’s perceived importance. In this paper, we introduce PCS as an auxiliary step following the enhancement process, aiming to refine the perceptual quality of the speech signal [28].
4 Experiments
4.1 Dataset
We utilized the VoiceBank-DEMAND dataset [30] for our study. This dataset comprises noisy speech recordings generated by mixing clean speech from the VoiceBank collection [36] with noise from the DEMAND dataset [37]. It encompasses 30 distinct speakers, with 28 speakers allocated for training and 2 for testing. Clean samples were mixed with noise samples at four signal-to-noise ratios (SNRs) during training ([0, 5, 10, 15] dB) and testing ([2.5, 7.5, 12.5, 17.5] dB). The training dataset comprised a total of 11,572 utterances, while the testing set included 824 utterances.
4.2 Experimental Setup
Following the practice suggested in [38], all recordings were downsampled from 48 kHz to 16 kHz. The evaluation metrics used to assess the SE performance include: (i) Wide-band PESQ  [23], (ii) Prediction of the signal distortion (CSIG), (iii) Prediction of the background intrusiveness (CBAK), (iv) Prediction of the overall speech quality (COVL), and (v) STOI  [24]. Higher values for these five metrics indicate better enhancement outcomes.
In our experimental setup, we systematically evaluate various aspects to comprehensively assess the performance of the proposed SEMamba system  [^1] The detailed results and model codes will be released at https://github.com/RoyChao19477/SEMamba.. First, we evaluate SEMamba’s performance using a basic model architecture, as depicted in Fig. 1. Next, we assess Mamba with a more advanced model architecture, as depicted in Fig. 2. Finally, we examine the compatibility of SEMamba with CL and PCS.
4.3 Experimental Results
Table 1 reports the experimental results with the basic SE architecture leveraging either the Mamba or Transformer block. The results show that Mamba delivers comparable or superior performance to the Transformer in both causal and non-causal configurations while utilizing fewer FLOPs and parameters than the Transformer.
Table 1: Comparison of basic SE architectures leveraging either Mamba or Transformer blocks.
Causal PESQ STOI FLOPs Para.
noisy-speech 1.97 0.92
Transformer Yes 2.76 0.94 2.26G 9.05M
Mamba Yes 2.76 0.94 0.76G 3.60M
Transformer No 2.84 0.94 2.26G 9.05M
Mamba No 2.85 0.94 1.06G 6.49M
Next, we compare Mamba and Transformer using a more advanced SE architecture, as shown in Fig. 2. The results are listed in Table 2. Please note that the original MP-SENet uses a Conformer as the core model. In our test experiments, Conformer and Transformer yield similar performance. All three SE systems presented in Table 2 are non-causal. Both Conformer and Mamba (Bi) utilized a bi-directional architecture, while Mamba (Uni) employed a uni-directional architecture.
By comparing the results in Tables 1 and 2, it is noted that Mamba achieves notably better performance when integrated with an advanced model architecture. Next, from Table 2, we observe that with a more advanced model architecture, Mamba yields better performance while requiring lower computational cost (in terms of FLOPs) than Conformer.
Table 2: Comparison between Mamba and Transformer with an advanced SE architecture (without CL and without PCS).
PESQ STOI FLOPs Para.
noisy-speech 1.97 0.92
Conformer 3.50 0.96 74.29G 2.05M
Mamba (Uni) 3.29 0.95 53.09G 1.41M
Mamba (Bi) 3.52 0.96 65.46G 2.25M
Finally, we compare SEMamba with several well-known SE methods on the VoiceBank-DEMAND dataset. The results are listed in Table 3. Note that three SE methods, namely S4DSE, S4ND-UNet, and Spiking-S4, derived based on the state-space model  [39, 40, 41] are included for comparison. We also list the results without (denoted as -CL) and with consistency loss, as well as with (denoted as +PCS) and without PCS, for a detailed performance comparison.
From Table 3, we first note that SEMamba (-CL) yields performance comparable to MP-SENet. The major distinction between SEMamba (-CL) and MP-SENet is that SEMamba (-CL) utilizes Mamba as the core model, whereas MP-SENet employs Conformer. The results again confirm that Mamba yields similar evaluation scores. Next, it is observed that SEMamba provides a high PESQ score of 3.55 when combining CL on the VoiceBank-DEMAND dataset. By additionally applying PCS, it can further improve to a score of 3.69, representing the SOTA results on this dataset.
Table 3: The results of SEMamba and several well-known SE solutions on the VoiceBank-DEMAND dataset.
Model PESQ CSIG CBAK COVL STOI
noisy-speech 1.97 0.92 3.34 2.63
SEGAN [38] 2.16 3.48 2.94 2.80
MetricGAN+ [26] 3.15 0.93 4.14 3.64
DPT [42] 3.33 0.96 4.58 4.00
CMGAN [17] 3.41 4.63 3.94 4.12 0.96
MP-SENet [18] 3.50 4.73 3.95 4.22 0.96
S4DSE [39] 2.55 3.94 3.00 3.23 0.93
S4ND-UNet [40] 3.15 4.52 3.62 3.85
Spiking-S4 [41] 3.39 4.92 2.64 4.31
SEMamba (-CL) 3.52 4.75 3.98 4.26 0.96
SEMamba 3.55 4.77 3.95 4.26 0.96
SEMamba (+PCS) 3.69 4.79 3.63 4.37 0.96
5 Conclusion
This study explored Mamba, a novel SSM incorporating a selective mechanism, for tackling the SE task. Comprehensive investigations involving basic and advanced SE neural schemes and employing traditional signal-level and metric-oriented objective functions were carried out. Experimental results showed that Mamba holds significant promise in advancing SE performance. In fact, on the VoiceBank-DEMAND dataset, SEMamba with PCS achieved a SOTA PESQ score of 3.69. In future research, we aim to explore Mamba’s potential in other speech generation tasks.